# Critical wetting in the (2+1)D solid-on-solid model

### v1

@@ 1. Introduction
The Solid-On-Solid (SOS) model above a wall with an attractive pinning force of $\lambda$ along the wall, is the following probability distribution over nonnegative height functions $\phi$ over $\Lambda_{n}=\{-\lfloor\frac{n}{2}\rfloor,\ldots,\lfloor\frac{n}{2}\rfloor\}^{2}$:
@@
A standard setting for the model sets 00 boundary condition on $\mathbb{Z}^{2}\setminus\Lambda$, which is incorporated into (1.1) by viewing $\phi$ as extended onto $\Lambda_{n}$ together with its outer vertex boundary $\partial\Lambda_{n}$, but forced to be 00 on $\partial\Lambda_{n}$.
@@
When $\lambda>0$, the pinning potential competes with the entropic repulsion; Chalker [5] showed that for all large $\beta$, there is a critical $\lambda_{w}(\beta)$ separating a localized regime of $\lambda>\lambda_{w}$ in which the interface height at the origin is tight: $\phi_{o}=O_{\textsc{p}}(1)$ as $n$ grows (and $\phi_{x}=0$ for most $x\in\Lambda_{n}$), and a delocalized wetting regime of $\lambda<\lambda_{w}$ where $\phi_{o}\to\infty$ (as does $\phi_{x}$ for almost all $x\in\Lambda_{n}$). Since that work, the two regimes have been investigated in more detail, e.g., [11, 1], culminating in works of Lacoin [9, 10]. These last two works identified
@@
The following observation of Lacoin [9] relating $\max\{0,\varphi\}$ for $\varphi\sim\widetilde{\mathbb{P}}$ to $\phi\sim\mathbb{P}$ will be very useful to us.
Observation 1.3.
@@
$\mathbb{P}(\phi)=\sum_{\varphi:\max\{0,\varphi\}=\phi}\widetilde{\mathbb{P}}(\varphi)\,.$
@@
In fact [9] gave a more general version of this equivalence, but the case $\lambda=\lambda_{w}$ is especially simple due to the equivalence between the ability in $\varphi$ for a singleton at height $\leq 0$ to take a geometric random variable with mean $e^{-4\beta}$ as a negative height, and the token of size $\frac{1}{1-e^{-4\beta}}$ that $\phi$ collects by being at height zero.
@@ Acknowledgements
The research of R.G. is supported in part by NSF DMS-2246780. The research of E.L. was supported by the NSF grant DMS-2054833.
@@ 2.1. A lifting map to gain entropy
$\partial_{e}V:=\{[u,v]:u\sim v,u\notin V,v\in V\}\,.$
Moreover, define $|\Delta S^{\prime}|$ as the change in disagreements attributed to the edge boundary of $S^{\prime}$:
$|\Delta S^{\prime}|=|\{u\sim v:u\in S^{\prime},v\notin S^{\prime},\varphi_{u}\geq\varphi_{v}\}|-|\{u\sim v:u\in S^{\prime},v\notin S^{\prime},\varphi_{u}<\varphi_{v}\}|\,.$ (2.1)
@@
$\sum_{A\subseteq\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W}\widetilde{\mathbb{P}}_{S}^{h}(U_{A}\varphi)\geq\widetilde{\mathbb{P}}_{S}^{h}(\varphi)e^{-\beta|\Delta S^{\prime}|-\beta|\partial_{e}W|-\lambda|W|}(1+\frac{1}{2}e^{-6\beta})^{|\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W|/5}\,.$
@@
~$\widetilde{\mathbb{P}}_{S}^{h}(U_{A}\varphi)\geq\widetilde{\mathbb{P}}_{S}^{h}(\varphi)\exp{(-\beta|\Delta S^{\prime}|-\beta|\partial_{e}A|-\beta|\partial_{e}W|-\lambda(|\mathsf{Q}_{\geq 2}(\varphi)|-|\mathsf{Q}_{\geq 2}(U_{A}\varphi)|))}\,.$ (2.2)
@@
$\sum_{A\subseteq\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W}\widetilde{\mathbb{P}}_{S}^{h}(U_{A}\varphi)\geq\widetilde{\mathbb{P}}_{S}^{h}(\varphi)e^{-\beta|\Delta S^{\prime}|-\beta|\partial_{e}W|-\lambda|W|}\prod_{i=1}^{|\mathcal{T}|}e^{-\lambda|T_{i}|}\sum_{A_{i}\subseteq T_{i}}e^{-\beta|\partial_{e}A_{i}|+\lambda|\mathsf{Q}_{\geq 2}(A_{i})|}\,,$ (2.4)
@@
$\sum_{A\subseteq\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W}\widetilde{\mathbb{P}}_{S}^{h}(U_{A}\varphi)\geq\widetilde{\mathbb{P}}_{S}^{h}(\varphi)e^{-\beta|\Delta S^{\prime}|-\beta|\partial_{e}W|-\lambda|W|}(1+\frac{1}{2}e^{-6\beta})^{|\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W|/5}\,.\qed$
@@
When $S=S^{\prime}=\Lambda_{n}$, $h=0$, and $W=\emptyset$, the above corresponds to the result of [6, Thm 1.2] that $\mathsf{Q}_{\geq 2}(\varphi)\leq C_{\beta}n$ (e.g., this bound holds for the choice $C_{\beta}=e^{7\beta}$). To make the present paper self-contained ($|\mathsf{Q}_{\geq 2}(\varphi)|=O(n)$ will be needed in Section 3), we show how to derive said bound as a corollary of the above.
@@ 2.2. Identifying the rate for downward tooth-like oscillations
For every simply connected $S$, every $h\geq 1$, and every pair of adjacent sites $x,y\in S$,
@@
~$\widetilde{\mathbb{P}}_{S}^{h}(|\mathcal{C}_{x,y}(\varphi)|=6,\varphi_{x}\leq 1,\varphi_{y}\leq 1,\varphi_{x}=0\text{ or }\varphi_{y}=0)\leq e^{-6\beta(h-1)}.$
@@ 3. Lower bound
Fix $k\geq 2$, and let $\mathcal{B}_{k}$ be the bad event that $|X_{k}(\varphi)|\geq e^{8\beta-5\beta k}n^{2}$, and $|\mathsf{Q}_{\geq 2}(\varphi)|\leq e^{7\beta}n$. Let $X^{\prime}_{k}(\varphi)$ be an arbitrary subset of disjoint ordered pairs of adjacent sites $(x,y)$ among $X_{k}(\varphi)$, noting that we may collect at least $|X_{k}(\varphi)|/7$ such pairs greedily (e.g., by listing the edges $xy$ of $X_{k}(\varphi)$, ordered $x<y$ lexicographically, proceeding sequentially and collecting each one that is not sharing a vertex with a previously selected edge).
@@
Under the identification $\phi_{v}=\max\{0,\varphi_{v}\}$ for all $v$, the sets of sites not at height $\{h_{n}^{*}-1,h_{n}^{*}\}$ are fixed. Thus, by Observation 1.3,

### v2

@@ 1. Introduction
The Solid-On-Solid (SOS) model above a wall with an attractive pinning force of $\lambda>0$ along the wall, is the following probability distribution over nonnegative height functions $\phi$ over $\Lambda_{n}=\{-\lfloor\frac{n}{2}\rfloor,\ldots,\lfloor\frac{n}{2}\rfloor\}^{2}$:
@@
A standard setting for the model sets 00 boundary condition on $\mathbb{Z}^{2}\setminus\Lambda_{n}$, which is incorporated into (1.1) by viewing $\phi$ as extended onto $\Lambda_{n}$ together with its outer vertex boundary $\partial\Lambda_{n}$, but forced to be 00 on $\partial\Lambda_{n}$.
@@
When $\lambda>0$, the pinning potential competes with the entropic repulsion; Chalker [5] showed that for all large $\beta$, there is a critical $\lambda_{w}(\beta)$ separating a localized regime of $\lambda>\lambda_{w}$ in which the interface height at the origin is tight: $\phi_{o}=O_{\textsc{p}}(1)$ as $n$ grows (and $\phi_{x}$ is a certain constant $k(\lambda)$ for most $x\in\Lambda_{n}$), and a delocalized wetting regime of $\lambda<\lambda_{w}$ where $\phi_{o}\to\infty$ (as does $\phi_{x}$ for almost all $x\in\Lambda_{n}$). Since that work, the two regimes have been investigated in more detail, e.g., [11, 1], culminating in works of Lacoin [9, 10]. These last two works identified
@@
Remark 1.3.
Lemma 2.6 in fact shows that all connected sets of vertices of height at least $h_{n}^{*}+1$ have exponential tails on their sizes, and by a union bound the largest one is $O(\log n)$ with high probability. One would expect, in analogy with the $\lambda=0$ case [4], that there is a unique macroscopic connected component of the $\{x:\phi_{x}\in\{h_{n}^{*}-1,h_{n}^{*}\}\}$, occupying all but some $\varepsilon_{\beta}$ fraction of $\Lambda_{n}$, whose scaling limit is a Wulff shape.
@@
The following observation of [6] relating $\max\{0,\varphi\}$ for $\varphi\sim\widetilde{\mathbb{P}}$ to $\phi\sim\mathbb{P}$ will be very useful.
Observation 1.4.
@@
$\mathbb{P}(\phi)=\sum_{\varphi:\max\{0,\varphi\}=\phi}\widetilde{\mathbb{P}}(\varphi)\,,$
@@
Lemma 3.1 of  [9] also used a relaxation of the floor constraint, but that one was more complicated and for general $\lambda$. The case $\lambda=\lambda_{w}$ with the state space $\widetilde{\Omega}$ as suggested in [6, Section 4] is especially clean due to the equivalence between the ability in $\varphi$ for a singleton at height $\leq 0$ to take a geometric random variable with mean $e^{-4\beta}$ as a negative height, and the token of size $\frac{1}{1-e^{-4\beta}}$ that $\phi$ collects by being at height zero.
@@ Acknowledgements
The authors thank the anonymous referee for helpful comments. The research of R.G. is supported in part by NSF DMS-2246780. The research of E.L. was supported by the NSF grant DMS-2054833.
@@ 2.1. A lifting map to gain entropy
$\partial_{e}V:=\{\{u,v\}:u\sim v,u\notin V,v\in V\}\,.$
Moreover, for a $\varphi$, define $\Delta S^{\prime}$ as the change in disagreements attributed to the edge boundary of $S^{\prime}$:
$\Delta S^{\prime}=\Delta S^{\prime}(\varphi):=|\{u\sim v:u\in S^{\prime},v\notin S^{\prime},\varphi_{u}\geq\varphi_{v}\}|-|\{u\sim v:u\in S^{\prime},v\notin S^{\prime},\varphi_{u}<\varphi_{v}\}|\,.$ (2.1)
@@
$\sum_{A\subseteq\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W}\widetilde{\mathbb{P}}_{S}^{h}(U_{A}\varphi)\geq\widetilde{\mathbb{P}}_{S}^{h}(\varphi)e^{-\beta\Delta S^{\prime}-\beta|\partial_{e}W|-\lambda|W|}(1+\frac{1}{2}e^{-6\beta})^{|\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W|/5}\,.$
@@
~$\widetilde{\mathbb{P}}_{S}^{h}(U_{A}\varphi)\geq\widetilde{\mathbb{P}}_{S}^{h}(\varphi)\exp{(-\beta\Delta S^{\prime}-\beta|\partial_{e}A|-\beta|\partial_{e}W|-\lambda(|\mathsf{Q}_{\geq 2}(\varphi)|-|\mathsf{Q}_{\geq 2}(U_{A}\varphi)|))}\,.$ (2.2)
@@
$\sum_{A\subseteq\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W}\widetilde{\mathbb{P}}_{S}^{h}(U_{A}\varphi)\geq\widetilde{\mathbb{P}}_{S}^{h}(\varphi)e^{-\beta\Delta S^{\prime}-\beta|\partial_{e}W|-\lambda|W|}\prod_{i=1}^{|\mathcal{T}|}e^{-\lambda|T_{i}|}\sum_{A_{i}\subseteq T_{i}}e^{-\beta|\partial_{e}A_{i}|+\lambda|\mathsf{Q}_{\geq 2}(A_{i})|}\,,$ (2.4)
@@
$\sum_{A\subseteq\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W}\widetilde{\mathbb{P}}_{S}^{h}(U_{A}\varphi)\geq\widetilde{\mathbb{P}}_{S}^{h}(\varphi)e^{-\beta\Delta S^{\prime}-\beta|\partial_{e}W|-\lambda|W|}(1+\frac{1}{2}e^{-6\beta})^{|\mathsf{Q}_{\geq 2}(\varphi)\cap S^{\prime}\setminus W|/5}\,.\qed$
@@
When $S=S^{\prime}=\Lambda_{n}$, $h=0$, and $W=\emptyset$, Lemma 2.2 corresponds to the result of [6, Thm 1.2] that $\mathsf{Q}_{\geq 2}(\varphi)\leq C_{\beta}n$ (e.g., this bound holds for the choice $C_{\beta}=e^{7\beta}$). To make the present paper self-contained ($|\mathsf{Q}_{\geq 2}(\varphi)|=O(n)$ will be needed in Section 3), we show how to derive said bound as a corollary of the above.
@@ 2.2. Identifying the rate for downward tooth-like oscillations
There exists $\varepsilon_{\beta}>0$ (going to zero as $\beta\to\infty$) such that for every simply connected $S$, every $h\geq 1$, and every pair of adjacent sites $x,y\in S$,
@@
~$\widetilde{\mathbb{P}}_{S}^{h}(|\mathcal{C}_{x,y}(\varphi)|=6,\varphi_{x}\leq 1,\varphi_{y}\leq 1,\varphi_{x}\neq 0\text{ or }\varphi_{y}\neq 0)\leq e^{-6\beta(h-1)}.$
@@ 3. Lower bound
Fix $k\geq 2$, and let $\mathcal{B}_{k}$ be the bad event that $|X_{k}(\varphi)|\geq e^{8\beta-5\beta k}n^{2}$, and $|\mathsf{Q}_{\geq 2}(\varphi)|\leq e^{7\beta}n$. (Note that the choice of $-5\beta k$ is somewhat arbitrary in that $5$ could be any number strictly smaller than $6$, since $e^{-6\beta k}n^{2}$ is, to first order, the expected number of sites at depth $k$ below $h_{n}^{*}$.)
Let $X^{\prime}_{k}(\varphi)$ be an arbitrary subset of disjoint ordered pairs of adjacent sites $(x,y)$ among $X_{k}(\varphi)$, noting that we may collect at least $|X_{k}(\varphi)|/7$ such pairs greedily (e.g., by listing the edges $xy$ of $X_{k}(\varphi)$, ordered $x<y$ lexicographically, proceeding sequentially and collecting each one that is not sharing a vertex with a previously selected edge).
@@
Under the identification $\phi_{v}=\max\{0,\varphi_{v}\}$ for all $v$, the sets of sites not at height $\{h_{n}^{*}-1,h_{n}^{*}\}$ are fixed. Thus, by Observation 1.4,